# From Hyperons to Hypernuclei: A New Route to Unravel Proton Spin Polarization

### v1

@@ Abstract
Ultra-relativistic nuclear collisions create the quark–gluon plasma (QGP) known as the hottest, least viscous, and most vortical fluid ever produced in terrestrial laboratories. Its vortical structure has been uncovered through the spin polarization of Lambda ($\Lambda$) hyperons, attributed to the spin–orbit coupling that transfers the system’s orbital angular momentum to the quark spin, which is then inherited by hadrons via quark recombination or coalescence. However, $\Lambda$ polarization reflects primarily the strange-quark component, leaving the spin dynamics of the up and down quarks largely unexplored. Although the proton is an ideal probe, its stability makes direct measurements experimentally challenging. Here, we propose to reveal proton spin polarization via hypertriton ($^{3}_{\Lambda}\text{H}$) measurements, exploiting the fact that spin information is preserved when polarized nucleons and $\Lambda$ coalesce to form hypertriton. We show that, over a broad range of collision energies, the polarizations of proton, $\Lambda$, and hypertriton are related by a simple linear scaling law. Since both $\Lambda$ and hypertriton polarizations can be measured via their self-analyzing weak decays, this linear relation provides a practical experimental avenue for accessing spin polarizations of protons and neutrons–the dominant baryonic degrees of freedom in nuclear collisions.
@@
Figure 1: Schematic illustration of decoding proton spin polarization in heavy-ion collisions. a, In non-central heavy-ion collisions, the overlap region of colliding nuclei carries a large orbital angular momentum $\vec{L}$, leading to spin polarization of quarks through the spin-orbit interaction. Hadrons emerging from the QGP at hadronization acquire spin polarization via coalescence of polarized quarks, and subsequently evolve through the hadronic phase until the kinetic freeze-out. b, Hypertritons ($^{3}_{\Lambda}\text{H}$) are polarized when formed via the coalescence of polarized nucleons and $\Lambda$ hyperons at the kinetic freeze-out. Both $^{3}_{\Lambda}\text{H}$ and $\Lambda$ undergo self-analyzing weak decays (e.g. $\Lambda\rightarrow\pi^{-}+p$ and $^{3}_{\Lambda}\text{H}$$\rightarrow\pi^{-}+^{3}$He), allowing their spin polarizations to be determined from the angular distributions ($dN/\sin\theta^{*}d\theta^{*}$) of their decay products. c, The proton spin polarization ($\mathcal{P}_{p}$) can then be reconstructed from the polarizations of $^{3}_{\Lambda}\text{H}$ and $\Lambda$ using a model-validated linear relation.
@@
In semi-central collisions, the QGP carries substantial orbital angular momentum, on the order of $10^{3}\hbar$, oriented perpendicular to the reaction plane [2]. Through the spin–orbit or spin–vorticity coupling, part of this angular momentum can be transferred to the spin degrees of freedom of quarks, and subsequently inherited by hadrons via quark coalescence or recombination at hadronization [3]. This mechanism gives rise to the phenomenon of global spin polarization [3, 4, 5, 6], analogous to the Barnett and Einstein–de Haas effect in condensed matter [7, 8]. Experiments at the Relativistic Heavy Ion Collider (RHIC) have observed global spin polarization of $\Lambda$ and $\bar{\Lambda}$ hyperons, confirming that the QGP is the most vortical fluid ever observed [2]. Since then, significant progress has been made to understand the role of spin degrees of freedom in the dynamics of QGP [9, 10, 11, 12, 13, 14, 15].
Despite these advances, a crucial piece in our understanding of spin dynamics in hot and dense QCD matter remains missing: the spin polarization of protons and neutrons—the fundamental constituents of visible matter and dominant baryonic degrees of freedom in nuclear collisions—has not been measured. Measuring proton spin polarization is of particular importance for addressing the polarization of light quarks in the QGP (beyond what $\Lambda$ polarization reveals about strange quarks) and the spin-dependent transport phenomenon [16, 17] such as the haronic spin Hall effect [18]. Unlike hyperons, however, these baryons do not decay within the detector volume, making conventional polarization analysis via self-analyzing weak decays infeasible. A direct measurement would require external polarimeters based on secondary interactions, which has not yet been implemented in current heavy-ion detectors [19].
@@
This relation is a direct reflection of the fact that in the spin wavefunction of the $^{3}_{\Lambda}\text{H}$ ($J^{\pi}=\frac{1}{2}^{+}$ [22]), the $\Lambda$ spin aligns with the total spin with 1/3 probability and anti-aligns with 2/3 probability. A schematic illustration of the proposed method for decoding proton spin polarization via $^{3}_{\Lambda}\text{H}$ production is shown in Fig. 1.
Figure 2: Collision energy dependence of global spin polarizations of $\Lambda$ hyperons, protons, and hypertritons. a, Global polarization of mid-rapidity ($|y|<1$) $\Lambda$ hyperons in Au+Au collisions at $\sqrt{s_{NN}}=7.7$–200 GeV and 20–50% centrality. The blue and orange bands represent model predictions from an extended AMPT model at hadronization and kinetic freeze-out stages, respectively, using an impact parameter $b=8$ fm. Experimental data for $\Lambda$ hyperons from the STAR collaboration are shown with error bars, where the uncertainties combine statistical and systematic contributions [2, 23]. b, Corresponding global polarization of proton at hadronization (blue band) and kinetic freeze-out (orange band). c, Same as in b, but for hypertritons. Uncertainties in theoretical results are statistical only.
@@
Figure 2 (b) depicts the beam energy dependence of proton spin polarization at the hadronization stage (blue band) and at kinetic freeze-out (orange band). At hadronization, the spin polarizations of nucleons and $\Lambda$ hyperons are found to be close to each other, owing to their similar masses and the same spin quantum numbers. Unlike the case for $\Lambda$ hyperons, whose polarization is significantly reduced during the hadronic phase, the nucleon polarization remains almost unchanged between hadronization and kinetic freeze-out stages. This difference results from a near cancellation between two competing effects on proton spin polarization: an enhancement from feed-down decays of primary $\Delta(1232)$ resonances (with spin-3/2), and a suppression from spin depolarization induced by inelastic scatterings with pions, particularly through the process $\pi N\rightarrow\Delta\rightarrow\pi N$.
According to the quark recombination or coalescence model, the $\Lambda$ hyperon’s polarization ($\mathcal{P}_{\Lambda}$) directly reflects that of the strange quark ($\mathcal{P}_{s}$), as the up-down ($ud$) diquark inside the $\Lambda$ forms a spin singlet and contributes no net spin, i.e., $\mathcal{P}_{\Lambda}\approx\mathcal{P}_{s}$ [3, 33]. Thus, the measured $\Lambda$ polarization provides little information on the spin polarization of $u$ and $d$ quarks in the QGP. In contrast, the proton spin polarization ($\mathcal{P}_{p}$) is directly related to the spin polarizations of $u$ and $d$ quarks and can be approximated as $\mathcal{P}_{p}\approx(4\mathcal{P}_{u}-\mathcal{P}_{d})/3$ [3]. Therefore, the experimental measurement of proton spin polarization would provide valuable insights into the spin dynamics of light quarks in the QGP.
@@
Figure 3: Decoding proton spin polarization in heavy-ion collisions. a, Collision-energy dependence of global proton spin polarization in non-central Au+Au collisions at $\sqrt{s_{NN}}=7.7$–200 GeV, evaluated at mid-rapidity ($|y|<1$). b, Transverse-momentum dependence of proton spin polarization at the energy of 7.7 GeV. The lower panels show the ratio of the extracted proton polarization to that expected from the vorticity-induced polarization. Coloured bands represent statistical uncertainties in theoretical results.
@@
where $\alpha_{H}$ (subscript $H$ stands for $\Lambda$ and $^{3}_{\Lambda}\text{H}$) is the decay parameter, the polarization $\mathcal{P}_{H}=(N_{H,\frac{1}{2}}-N_{H,-\frac{1}{2}})/(N_{H,\frac{1}{2}}+N_{H,-\frac{1}{2}})$ characterizes the imbalance between the number of mother particle ($N_{H,\pm\frac{1}{2}}$) in different spin states, and $\theta^{*}$ denotes the angle between the daughter’s momentum and the spin direction of the mother particle. These weak decays are self-analyzing, as the momenta of the daughters carry direct information about the mother’s polarization.
@@
Decoding proton spin polarization from $\Lambda$ and $^{3}_{\Lambda}$H polarizations We now demonstrate how to reveal the proton spin polarization from polarizations of hypertriton and $\Lambda$ hyperon. The spin polarization of $^{3}_{\Lambda}\text{H}$ inherits those of nucleons and $\Lambda$ hyperons at their kinetic freeze-out. This becomes evident if one assumes that the polarizations of $\Lambda$ hyperons and nucleons are independent of their momenta, in which case the polarization of $^{3}_{\Lambda}\text{H}$ from Eq. (From Hyperons to Hypernuclei: A New Route to Unravel Proton Spin Polarization) can be simplified as [21]
@@
where we have neglected higher-order terms (since polarizations are typically only a few percent) and assumed isospin symmetry, i.e., $\mathcal{P}_{p}\approx\mathcal{P}_{n}$. This relation explains why the $^{3}_{\Lambda}$H polarization changes only slightly from hadronization to kinetic freeze-out in Fig. 2 (c), in contrast to the larger change for $\Lambda$ polarization in Fig. 2 (a). Inverting Eq. (5) yields the expression for proton spin polarization given in Eq. (1).
@@
Figure 3 (b) further shows the $p_{T}$ dependence of $\mathcal{P}_{p}(p_{T})$ and $(3\mathcal{P}_{\text{${}^{3}_{\Lambda}\text{H}$}}(3p_{T})+\mathcal{P}_{\Lambda}(p_{T}))/4$ in Au+Au collisions at $\sqrt{s_{NN}}=7.7$ GeV. Both quantities exhibit a decreasing trend with $p_{T}$ and remain in close agreement throughout. Their ratio, shown in the lower panel, stays near unity within uncertainties, further validating the effectiveness of Eq. (1) in capturing both the magnitude and momentum dependence of proton spin polarization.
@@
Conclusions The spin information of protons and neutrons-the dominant baryonic constituents-is largely inaccessible in nuclear collisions due to their stabilities. To enable such measurements, we have proposed a novel method to reveal the proton spin polarization through independent measurements of the polarizations of $\Lambda$ hyperons and $^{3}_{\Lambda}\text{H}$. To validate this method, we perform transport model calculations incorporating spin-dependent scatterings and resonance feed-down effects during the hadronic phase of Au+Au collisions at $\sqrt{s_{NN}}=7.7$–200 GeV. We find that the proton polarization is related to the measurable $\Lambda$ and $^{3}_{\Lambda}\mathrm{H}$ polarizations through a simple and robust linear relation given by $\mathcal{P}_{p}\approx\left(3\mathcal{P}_{\text{${}^{3}_{\Lambda}\text{H}$}}+\mathcal{P}_{\Lambda}\right)/4$.
Accessing the proton spin polarization with our proposed method would open a unique window into the spin dynamics of up and down quarks, complementing existing insights from $\Lambda$ polarization that probe strange quarks. It also offers opportunities to explore spin-dependent transport phenomena in hadronic matter [16, 17], constrain the equation of state of spin-polarized nuclear matter [43], and investigate the polarized parton distributions in nucleons [44, 45]. This method can be readily applied to ongoing and upcoming experiments at RHIC, the High Intensity heavy-ion Accelerator Facility (HIAF) [46], the Nuclotron-based Ion Collider Facility (NICA) [47], and the Facility for Antiproton and Ion Research (FAIR) [48].
@@
Acknowledgments The authors thank helpful discussions with Lie-Wen Chen, Zuo-Tang Liang, Shu Lin, Shi Pu, Yi-Feng Sun, Rui Wang, Jun Xu, and Zhen Zhang. This work was supported in part by the National Key Research and Development Project of China under Grant No. 2024YFA1612500 and No. 2024YFA1610802; the National Natural Science Foundation of China under contract No. 12422509, No. 12375121, No. 124B2102, No. 12147101, No. 12275054, No. 12025501, and No. 12061141008; the Natural Science Foundation of Shanghai under Grant No. 23JC1400200 and No. 23590780100; the Shanghai Pilot Program for Basic Research-Fudan University 21TQ1400100(22TQ006); the Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008; the STCSM under Grant No. 23590780100; and the U.S. Department of Energy under Award No. DE-SC0015266. The computations in this research were performed using the CFFF platform of Fudan University.
@@
where $n_{F}=1/[1+\exp(\beta\cdot p\mp\mu/T)]$ is the Fermi-Dirac distribution function for particles ()(-) and antiparticles $(+)$. In heavy-ion collisions, $n_{F}\approx 1$ due to the high temperature and low baryon and strangeness densities. In Eq. (7), the thermal vorticity tensor $\varpi_{\mu\nu}$ is defined as $\varpi_{\mu\nu}=-\frac{1}{2}(\partial_{\mu}\beta_{\nu}-\partial_{\nu}\beta_{\mu})$ with $\beta^{\mu}=u^{\mu}/T$ being the four-velocity (satisfying the normalization condition $u^{\mu}u_{\mu}=1$) of the fluid divided by temperature. The temperature field and velocity field are obtained by solving the eigen equation $T^{\mu\nu}u_{\mu}=\epsilon u^{\mu}$, where $T^{\mu\nu}$ is the energy-momentum tensor, $\epsilon$ denotes the energy density of fluid cells in its local rest frame [51]. The temperature field is estimated as $T=0.199~\text{GeV}\left(\frac{\epsilon/\gamma_{q}}{1+3N_{f}/4}\right)^{1/4}$ [52] with a fugacity factor $\gamma_{q}=\frac{1}{2}\left(\left(\frac{N_{+}}{N_{-}}\right)^{N_{f}/2}+\left(\frac{N_{-}}{N_{+}}\right)^{N_{f}/2}\right)$ where $N_{+}$ ($N_{-}$) represents the number of (anti)quarks in a cell summed over all events and $N_{f}=3$ corresponds to the number of quark flavors. Note that these fields are determined by averaging over a sufficiently large number (typically $10^{5}$) of events from AMPT and the spin polarization vector is calculated in the particle’s rest frame with a Lorentz boost [51].
@@
where $\mathcal{C}_{M\rightarrow D}$ is the spin transfer coefficient for the decay [53, 54, 55]. For strong decays such as $\Delta\to N\pi$, $\Sigma^{*}\to\Lambda\pi$, and $N^{*}\to N\pi$, the spin transfer coefficient is $\mathcal{C}_{N^{*}\rightarrow N}=1$. In contrast, for the electromagnetic decay $\Sigma^{0}\to\Lambda\gamma$, the spin transfer coefficient is negative, $\mathcal{C}_{\Sigma^{0}\rightarrow\Lambda}=-\tfrac{1}{3}$ [53]. As a result, $\Lambda$ polarization is enhanced by feed-down from $\Sigma^{*}$ decays but suppressed by contributions from $\Sigma^{0}$. Likewise, the feed-down from primary $\Delta$ resonances leads to an enhancement of nucleon polarization. For weak decays $\Xi\to\Lambda\pi$ and $\Omega\to\Lambda K$, which are not included in the present study, their effects on $\Lambda$ polarization are negligible because of the relative small number of $\Xi$ and $\Omega$ [56, 32].
@@
where $p^{\mu}$ is the four-momentum, $d\sigma_{\mu}$ represents the normal vector of the hypersurface of freeze-out particles, $\bar{f}({\bf x}_{i},{\bf p}_{i})$ denotes their spin-averaged phase-space distributions with the coordinate $x_{i}$ and momentum $p_{i}$ in the frame of the emission source, and $W_{\text{${}^{3}_{\Lambda}\text{H}$}}$ is the Wigner function of $^{3}_{\Lambda}\text{H}$, which we take to have a Gaussian form, $W_{{}^{3}_{\Lambda}\text{H}}=8^{2}\exp\left[-\frac{\rho^{2}}{\sigma_{\rho}^{2}}-\frac{\lambda^{2}}{\sigma_{\lambda}^{2}}-p_{\rho}^{2}\sigma_{\rho}^{2}-p_{\lambda}^{2}\sigma_{\lambda}^{2}\right]$ with $\rho$, $\lambda$, $p_{\rho}$, and $p_{\lambda}$ denoting the relative distance in coordinate and momentum spaces [58]. The parameters used in the Wigner function are $\sigma_{\rho}=2.26$ fm to reproduce the empirical deuteron root-mean-squared radius of $\sqrt{\left<r_{d}^{2}\right>}=1.96$ fm [39] and $\sigma_{\lambda}=6.52$ fm, which corresponds to the world average $\Lambda$ separation energy of $B_{\Lambda}=0.164$ MeV in $^{3}_{\Lambda}\text{H}$. Using the phase-space information of nucleons and $\Lambda$ hyperons at kinetic freeze-out (i.e., their positions and momenta at the last interactions) from the extended AMPT model, the integrated yield of hypertriton from Eq. (10) is simplified to Eq. (From Hyperons to Hypernuclei: A New Route to Unravel Proton Spin Polarization).

### v3

@@ Abstract
Ultra-relativistic nuclear collisions create the quark–gluon plasma (QGP) known as the hottest, least viscous, and most vortical fluid ever produced in terrestrial laboratories. Its vortical structure has been uncovered through the spin polarization of Lambda ($\Lambda$) hyperons, attributed to the spin–orbit coupling that transfers the system’s orbital angular momentum to the quark spin, which is then inherited by hadrons via quark recombination or coalescence. However, $\Lambda$ polarization reflects primarily the strange-quark component, leaving the spin dynamics of the up and down quarks largely unexplored. Although the proton is an ideal probe, its stability makes direct measurements experimentally challenging. Here, we propose to unravel proton spin polarization via hypertriton ($^{3}_{\Lambda}\text{H}$) measurements, exploiting the fact that spin information is preserved when polarized nucleons and $\Lambda$ coalesce to form hypertriton. We show that, over a broad range of collision energies, the polarizations of proton, $\Lambda$, and hypertriton are related by a simple linear scaling law. Since both $\Lambda$ and hypertriton polarizations can be measured via their self-analyzing weak decays, this linear relation provides a practical experimental avenue for accessing spin polarizations of protons and neutrons–the dominant baryonic degrees of freedom in nuclear collisions.
@@
Figure 1: Schematic illustration of decoding proton spin polarization in heavy-ion collisions. a, In non-central heavy-ion collisions, the overlap region of colliding nuclei carries a large orbital angular momentum $L$, leading to spin polarization of quarks through the spin-orbit interaction. Hadrons emerging from the QGP at hadronization acquire spin polarization via coalescence of polarized quarks, and subsequently evolve through the hadronic phase until the kinetic freeze-out. b, Hypertritons ($^{3}_{\Lambda}\text{H}$) are polarized when formed via the coalescence of polarized nucleons and $\Lambda$ hyperons at the kinetic freeze-out. Both $^{3}_{\Lambda}\text{H}$ and $\Lambda$ undergo self-analyzing weak decays (e.g. $\Lambda\rightarrow\pi^{-}+p$ and $^{3}_{\Lambda}\text{H}$$\rightarrow\pi^{-}+^{3}$He), allowing their spin polarizations to be determined from the angular distributions ($dN/\sin\theta^{*}d\theta^{*}$) of their decay products. c, The proton spin polarization ($\mathcal{P}_{p}$) can then be reconstructed from the polarizations of $^{3}_{\Lambda}\text{H}$ and $\Lambda$ using a model-validated linear relation.
@@
In semi-central collisions, the QGP carries substantial orbital angular momentum, on the order of $10^{3}\hbar$, oriented perpendicular to the reaction plane [2]. Through the spin–orbit or spin–vorticity coupling, part of this angular momentum can be transferred to the spin degrees of freedom of quarks, and subsequently inherited by hadrons via quark coalescence or recombination at hadronization [3]. This mechanism gives rise to the phenomenon of global spin polarization [3, 4, 5, 6], analogous to the Barnett and Einstein–de Haas effect in condensed matter [7, 8]. Experiments at the Relativistic Heavy Ion Collider (RHIC) have observed global spin polarization of $\Lambda$ and $\bar{\Lambda}$ hyperons, confirming that the QGP is the most vortical fluid ever observed [2]. Since then, significant progress has been made to understand the role of spin degrees of freedom in the dynamics of QGP [9, 10, 11, 12, 13, 14, 15, 16].
Despite these advances, a crucial piece in our understanding of spin dynamics in hot and dense QCD matter remains missing: the spin polarization of protons and neutrons—the fundamental constituents of visible matter and dominant baryonic degrees of freedom in nuclear collisions—has not been measured. Measuring proton spin polarization is of particular importance for addressing the polarization of light quarks in the QGP (beyond what $\Lambda$ polarization reveals about strange quarks) and the spin-dependent transport phenomenon [17, 18] such as the hadronic spin Hall effect [19]. Unlike hyperons, however, these baryons do not decay within the detector volume, making conventional polarization analysis via self-analyzing weak decays infeasible. A direct measurement would require external polarimeters based on secondary interactions, which has not yet been implemented in current heavy-ion detectors [20].
@@
This relation is a direct reflection of the fact that in the spin wavefunction of the $^{3}_{\Lambda}\text{H}$ ($J^{\pi}=\frac{1}{2}^{+}$ [23]), the $\Lambda$ spin aligns with the total spin with 1/3 probability and anti-aligns with 2/3 probability. A schematic illustration of the proposed method for decoding proton spin polarization via $\Lambda$ and $^{3}_{\Lambda}\text{H}$ production is shown in Fig. 1.
Figure 2: Collision energy dependence of global spin polarizations of $\Lambda$ hyperons, protons, and hypertritons. a, Global polarization of mid-rapidity ($|y|<1$) $\Lambda$ hyperons in Au+Au collisions at $\sqrt{s_{NN}}=7.7$–200 GeV and 20–50% centrality. The blue and orange bands represent model predictions from an extended AMPT model at hadronization and kinetic freeze-out stages, respectively, using an impact parameter $b=8$ fm. Experimental data for $\Lambda$ hyperons from the STAR collaboration are shown with error bars, where the uncertainties combine statistical and systematic contributions [2, 24]. b, Corresponding global polarization of proton at hadronization (blue band) and kinetic freeze-out (orange band). c, Same as in b, but for hypertritons. Uncertainties in theoretical results are from statistics.
@@
Figure 2 (b) depicts the beam energy dependence of proton spin polarization at the hadronization stage (blue band) and at kinetic freeze-out (orange band). At hadronization, the spin polarizations of nucleons and $\Lambda$ hyperons are found to be close to each other, owing to their similar masses and the same spin quantum numbers. Unlike the case for $\Lambda$ hyperons, whose polarization is significantly reduced during the hadronic phase, the nucleon polarization remains almost unchanged between hadronization and kinetic freeze-out stages. This difference results from a near cancellation between two competing effects on proton spin polarization: an enhancement from feed-down decays of primary $\Delta(1232)$ resonances (with spin-3/2), and a suppression from spin depolarization induced by quasi-elastic scatterings with pions, particularly through the process $\pi N\rightarrow\Delta\rightarrow\pi N$.
According to the quark recombination or coalescence model, the $\Lambda$ hyperon’s polarization ($\mathcal{P}_{\Lambda}$) directly reflects that of the strange quark ($\mathcal{P}_{s}$), as the up-down ($ud$) diquark inside the $\Lambda$ forms a spin singlet and contributes no net spin, i.e., $\mathcal{P}_{\Lambda}\approx\mathcal{P}_{s}$ [3, 34]. Thus, the measured $\Lambda$ polarization provides little information on the spin polarization of $u$ and $d$ quarks in the QGP. In contrast, the proton spin polarization ($\mathcal{P}_{p}$) is directly related to the spin polarizations of $u$ and $d$ quarks and can be approximated in the SU(6) quark model as $\mathcal{P}_{p}\approx(4\mathcal{P}_{u}-\mathcal{P}_{d})/3$  [3]. Therefore, the experimental measurement of proton spin polarization would provide valuable insights into the spin dynamics of light quarks in the QGP. Moreover, it would also shed light on the long-standing proton spin puzzle [35], given that first-principles Lattice QCD calculations indicate that quarks account for only about 38% of the proton’s intrinsic spin [36].
@@
Figure 3: Decoding proton spin polarization in heavy-ion collisions. a, Collision-energy dependence of global proton spin polarization in non-central Au+Au collisions at $\sqrt{s_{NN}}=7.7$–200 GeV, evaluated at mid-rapidity ($|y|<1$). b, Transverse-momentum dependence of proton spin polarization at the energy of 7.7 GeV. The lower panels show the ratio of the extracted value of proton polarization (orange bands) to its true value (blue bands) in model calculations. Coloured bands represent statistical uncertainties in theoretical results.
@@
where $\alpha_{H}$ (subscript $H$ stands for $\Lambda$ and $^{3}_{\Lambda}\text{H}$) is the decay parameter, the polarization $\mathcal{P}_{H}=(N_{H,\frac{1}{2}}-N_{H,-\frac{1}{2}})/(N_{H,\frac{1}{2}}+N_{H,-\frac{1}{2}})$ characterizes the imbalance between the number of mother particle ($N_{H,\pm\frac{1}{2}}$) in different spin states, and $\theta^{*}$ denotes the angle between the daughter’s momentum and the spin direction of the mother particle. The decay parameter of $^{3}_{\Lambda}\text{H}$ is linearly proportional to that of $\Lambda$, i.e., $\alpha_{{}^{3}_{\Lambda}\text{H}}\approx-\alpha_{\Lambda}/2.58$ [22]. These decays are self-analyzing, as the momenta of the daughters carry direct information about the mother’s polarization.
@@
Decoding proton spin polarization from $\Lambda$ and $^{3}_{\Lambda}$H polarizations We now demonstrate how to unravel the proton spin polarization from polarizations of hypertriton and $\Lambda$ hyperon. The spin polarization of $^{3}_{\Lambda}\text{H}$ inherits those of nucleons and $\Lambda$ hyperons at their kinetic freeze-out. This becomes evident if one assumes that the polarizations of $\Lambda$ hyperons and nucleons are independent of their momenta, in which case the polarization of $^{3}_{\Lambda}\text{H}$ from Eq. (From Hyperons to Hypernuclei: A New Route to Unravel Proton Spin Polarization) can be simplified as [22]
@@
where we have neglected higher-order terms (since polarizations are typically only a few percent) and assumed isospin symmetry, i.e., $\mathcal{P}_{p}\approx\mathcal{P}_{n}$. The dominance of proton polarization in Eq. (5) explains why the $^{3}_{\Lambda}$H polarization changes only slightly from hadronization to kinetic freeze-out in Fig. 2 (c), in contrast to the larger change for $\Lambda$ polarization in Fig. 2 (a). Inverting Eq. (5) yields the expression for proton spin polarization given in Eq. (1).
@@
Figure 3 (b) further shows the $p_{T}$ dependence of $\mathcal{P}_{p}(p_{T})$ and $(3\mathcal{P}_{\text{${}^{3}_{\Lambda}\text{H}$}}(3p_{T})+\mathcal{P}_{\Lambda}(p_{T}))/4$ in Au+Au collisions at $\sqrt{s_{NN}}=7.7$ GeV. Both quantities exhibit a decreasing trend with $p_{T}$ and remain in close agreement throughout. Their ratio, shown in the lower panel, stays near unity within uncertainties, further validating the effectiveness of Eq. (1) in capturing both the magnitude and momentum dependence of proton spin polarization. The same agreements hold for other collision energies.
@@
Conclusions The spin information of protons and neutrons-the dominant baryonic constituents in nuclear collisions-is largely inaccessible in nuclear collisions due to their stabilities. To enable such measurements, we have proposed a new route to unravel the proton spin polarization through independent measurements of the polarizations of $\Lambda$ hyperons and $^{3}_{\Lambda}\text{H}$. To validate this method, we perform transport model calculations incorporating spin-dependent scatterings and resonance feed-down effects during the hadronic phase of Au+Au collisions at $\sqrt{s_{NN}}=7.7$–200 GeV. We find that the proton polarization is related to the measurable $\Lambda$ and $^{3}_{\Lambda}\mathrm{H}$ polarizations through a simple and robust linear relation given by $\mathcal{P}_{p}\approx\left(3\mathcal{P}_{\text{${}^{3}_{\Lambda}\text{H}$}}+\mathcal{P}_{\Lambda}\right)/4$.
Accessing the proton spin polarization with our proposed method would opens a new window on the spin dynamics of up and down quarks, complementing existing insights from $\Lambda$ polarization that primarily probes strange quarks. It further enables studies of spin-dependent transport phenomena in hadronic matter [17, 18], provides constraints on the equation of state of spin-polarized nuclear matter [46], and helps resolving the proton spin puzzle in QCD [35, 47]. Importantly, this method can be readily applied to ongoing and upcoming experiments at RHIC, the High Intensity heavy-ion Accelerator Facility (HIAF) [48], the Nuclotron-based Ion Collider Facility (NICA) [49], and the Facility for Antiproton and Ion Research (FAIR) [50].
@@
Acknowledgments The authors thank helpful discussions with Lie-Wen Chen, Zuo-Tang Liang, Shu Lin, Xiao-Feng Luo, Guo-Liang Ma, Shi Pu, Yi-Feng Sun, Rui Wang, Jun Xu, Zhen Zhang, and Yong Zhao. This work was supported in part by the National Key Research and Development Project of China under Grant No. 2024YFA1612500 and No. 2024YFA1610802; the National Natural Science Foundation of China under contract No. 12422509, No. 12375121, No. 124B2102, No. 12147101, No. 12275054, No. 12025501, and No. 12061141008; the Natural Science Foundation of Shanghai under Grant No. 23JC1400200 and No. 23590780100; the Shanghai Pilot Program for Basic Research-Fudan University 21TQ1400100(22TQ006); the Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008; the STCSM under Grant No. 23590780100; and the U.S. Department of Energy under Award No. DE-SC0015266. The computations in this research were performed using the CFFF platform of Fudan University.
@@
where $n_{F}=1/[1+\exp(\beta\cdot p\mp\mu/T)]$ is the Fermi-Dirac distribution function for particles ()(-) and antiparticles $(+)$. In heavy-ion collisions, $n_{F}\approx 1$ due to the high temperature and low baryon and strangeness densities. In Eq. (7), the thermal vorticity tensor $\varpi_{\mu\nu}$ is defined as $\varpi_{\mu\nu}=-\frac{1}{2}(\partial_{\mu}\beta_{\nu}-\partial_{\nu}\beta_{\mu})$ with $\beta^{\mu}=u^{\mu}/T$ being the four-velocity (satisfying the normalization condition $u^{\mu}u_{\mu}=1$) of the fluid divided by temperature. The temperature field and velocity field are obtained by solving the eigen equation $T^{\mu\nu}u_{\mu}=\epsilon u^{\mu}$, where $T^{\mu\nu}$ is the energy-momentum tensor and $\epsilon$ denotes the energy density of fluid cells in its local rest frame [53]. The temperature field is estimated as $T=0.199~\text{GeV}\left(\frac{\epsilon/\gamma_{q}}{1+3N_{f}/4}\right)^{1/4}$ [54] with a fugacity factor $\gamma_{q}=\frac{1}{2}\left(\left(\frac{N_{+}}{N_{-}}\right)^{3/2}+\left(\frac{N_{-}}{N_{+}}\right)^{3/2}\right)$ where $N_{+}$ ($N_{-}$) represents the number of (anti)quarks in a cell summed over all events. Note that these fields are determined by averaging over a sufficiently large number (typically $10^{5}$) of events from AMPT and the spin polarization vector is calculated in the particle’s rest frame with a Lorentz boost [53].
@@
where $\mathcal{C}_{M\rightarrow D}$ is the spin transfer coefficient for the decay [55, 56, 57]. For strong decays such as $\Delta\to N\pi$, $\Sigma^{*}\to\Lambda\pi$, and $N^{*}\to N\pi$, the spin transfer coefficient is $\mathcal{C}_{M\rightarrow D}=1$. In contrast, for the electromagnetic decay $\Sigma^{0}\to\Lambda\gamma$, the spin transfer coefficient is negative, $\mathcal{C}_{\Sigma^{0}\rightarrow\Lambda}=-\tfrac{1}{3}$ [55]. As a result, $\Lambda$ polarization is enhanced by feed-down from $\Sigma^{*}$ decays but suppressed by contributions from $\Sigma^{0}$. Likewise, the feed-down from primary $\Delta$ resonances leads to an enhancement of nucleon polarization. For weak decays $\Xi\to\Lambda\pi$ and $\Omega\to\Lambda K$, which are not included in the present study, their effects on $\Lambda$ polarization are negligible because of the relative small number of $\Xi$ and $\Omega$ [58, 33].
@@
where $p^{\mu}$ is the four-momentum, $d\sigma_{\mu}$ represents the normal vector of the hypersurface of freeze-out particles, $\bar{f}({\bf x}_{i},{\bf p}_{i})$ denotes their spin-averaged phase-space distributions with the coordinate ${\bf x}_{i}$ and momentum ${\bf p}_{i}$ in the frame of the emission source, and $W_{\text{${}^{3}_{\Lambda}\text{H}$}}$ is the Wigner function of $^{3}_{\Lambda}\text{H}$, which we take to have a Gaussian form, $W_{{}^{3}_{\Lambda}\text{H}}=8^{2}\exp\left[-\frac{\rho^{2}}{\sigma_{\rho}^{2}}-\frac{\lambda^{2}}{\sigma_{\lambda}^{2}}-p_{\rho}^{2}\sigma_{\rho}^{2}-p_{\lambda}^{2}\sigma_{\lambda}^{2}\right]$ with $\rho$, $\lambda$, $p_{\rho}$, and $p_{\lambda}$ denoting the relative distance in coordinate and momentum spaces [60]. The parameters used in the Wigner function are $\sigma_{\rho}=2.26$ fm to reproduce the empirical deuteron root-mean-squared radius of $\sqrt{\left<r_{d}^{2}\right>}=1.96$ fm [42] and $\sigma_{\lambda}=6.52$ fm, which corresponds to the world average $\Lambda$ separation energy of $B_{\Lambda}=0.164$ MeV in $^{3}_{\Lambda}\text{H}$. Using the phase-space information of nucleons and $\Lambda$ hyperons at kinetic freeze-out (i.e., their positions and momenta at the last interactions) from the extended AMPT model, the integrated yield of hypertriton from Eq. (10) is simplified to Eq. (From Hyperons to Hypernuclei: A New Route to Unravel Proton Spin Polarization).